The mercury thread of a clinical thermometer reads 20.5cm at ice point and 35.7 point at steam point what will be it's temperature when it's reads 65 Celsius on the Celsius scale
1
Expert's answer
2018-01-24T09:40:08-0500
We can use the formula for calculating the temperature on an unmarked thermometer:
θ=(x-x_0)/(x_100-x_0 ) 100℃
In our case:
65℃=(x-20.5)/(35.7-20.5) 100℃
0.65=(x-20.5)/15.2
x=20.5+0.65(15.2)=30.4 cm.